# Neodymium ions as charge reservoir in NdNiO2: from lack of long range order to electron-doping-induced antiferromagnetism

### v1

@@ Abstract
We study magnetism in the electron-doped infinite-layer nickelate NdNiO2. We perform an unrestricted Hartree-Fock calculation for a tight-binding model which contains both nickel and neodymium orbitals. We reproduce the self-doping effect, which is the escape of charge onto the neodymium bands. By fixing all free parameters to realistic values we find that undoped NdNiO2 lies right outside the antiferromagnetic (AFM) region of the phase diagram. This is consistent with experiments, which find no long-range order in the ground state of NdNiO2, yet see short-range AFM correlations and broad magnetic excitations. Finally, we find that the self-doping effect leads to a dramatic increase in the stability of the AFM solution upon electron doping – a behavior that is strikingly different from what is, for instance, observed in the cuprates.
@@ II Model and Methods
$H_{t}=\sum_{ij}K_{i,j}=\sum_{ij}T(\mathbf{r}(i,j))\otimes k_{i,j}\otimes\sigma_{0}$ (3)
where $K_{i,j}$ is the hopping matrix between the unit cells $j$ and $i$, which can be written as a product of: (i) matrix $T(\mathbf{r}(i,j))$, the translation by the vector $\mathbf{r}(i,j)$, pointing from unit cell $j$ to unit cell $i$, with the lattice constant $a=1$; (ii) matrices $k_{i,j}$, each of dimension three, describing hopping elements in the orbital sector; (iii) matrix $\sigma_{0}$, identity in the spin sector. In reciprocal space this becomes
$H_{t}=\sum_{q}\sum_{r(i,j)}e^{i\mathbf{r}(i,j)\cdot\mathbf{q}}P_{q}\otimes k_{i,j}\otimes\sigma_{0},$ (4)
where $P_{q}$ is the projection onto the pseudomomentum $q$. The kinetic energy is quite rich, for it follows a realistic tight-binding model with hopping elements up to third neighbor [32] and therefore we need nine matrices $k_{i,j}$ to describe it. We provide those in Appendix A, Eq. (30). $k_{i,j}$ entries are all given in eV and follow [32]; matrices not given in (30) are zero.
@@
$\mathcal{H}^{\prime}=U\sum_{j}n_{j3\uparrow}n_{j3\downarrow},$ (5)
@@
$n_{j\sigma}=c^{\dagger}_{j3\sigma}c_{j3\sigma}.$ (6)
We treat $\mathcal{H}^{\prime}$ (5) in two ways. We use unrestricted Hartree-Fock (UHF) and restricted Hartree-Fock (RHF). We assume $U=4.5$ eV. Note that this value gives the $U/t\approx 12$ (where $t$ is the Ni-Ni nearest neighbor hopping element), basically in agreement with the recently postulated ratio used to describe the RIXS experiments on nickelates [33].
@@ II.1 Unrestricted Hartree-Fock approximation
In the above, $P_{j}$ is the projector onto the $j$-th unit cell, $P_{3}$ is the projector onto the nickel $d_{x^{2}-y^{2}}$ orbital, $\sigma_{0}$ is identity in the spin space and $\sigma_{\alpha},\;\alpha=x,y,z$ are the standard Pauli matrices. Furthermore,
$\begin{array}[]{c}\rho_{j\sigma}=\langle n_{j\sigma}\rangle\\ \Delta_{j}=\langle c^{\dagger}_{j\uparrow}c_{j\downarrow}\rangle,\end{array}$ (10)
[…]
$\Delta E_{UHF}=U\sum_{j}\left(-\rho_{j\uparrow}\rho_{j\downarrow}+|\Delta_{j}|^{2}\right).$ (11)
@@
In the RHF approach, we assume $\Delta_{j}=0$. We plug this into (8) and go to reciprocal space
$\begin{array}[]{c}\rho_{q\sigma}=\sum_{j}e^{i\mathbf{q}\cdot\mathbf{r}(j,0)}\rho_{j\sigma}\\ c_{q\sigma}=\frac{1}{\sqrt{N}}\sum_{j}e^{i\mathbf{q}\cdot\mathbf{r}(j,0)}c_{j\sigma},\end{array}$ (12)
where $N$ is the number of unit cells. Furthermore, we put restriction on $\rho_{q\sigma}$, limiting ourselves to ferromagnetic (FM) and antiferromagnetic (AFM) instabilities, namely
$\begin{array}[]{c}\rho_{q\uparrow}=\rho_{0\uparrow}\delta(\mathbf{q})+\rho_{1}\delta(\mathbf{q}-(\pi,\pi))\\ \rho_{q\downarrow}=\rho_{0\downarrow}\delta(\mathbf{q})-\rho_{1}\delta(\mathbf{q}-(\pi,\pi)).\end{array}$ (13)
@@
$\begin{split}H^{\prime}_{RHF}&=U\sum_{q}\Big{\{}P_{q}\otimes P_{3}\\ &\otimes\left[\frac{1}{2}\rho_{0\downarrow}\left(\sigma_{0}+\sigma_{z}\right)+\frac{1}{2}\rho_{0\uparrow}\left(\sigma_{0}-\sigma_{z}\right)\right]\\ &+T((\pi,\pi))\otimes P_{3}\otimes\rho_{1}\sigma_{z}\Big{\}}.\end{split}$ (16)
Above, $P_{q}$ is the projection onto the pseudomomentum $q$ and $T((\pi,\pi))$ is the translation in reciprocal space by the vector $(\pi,\pi)$.
@@
$\begin{split}\Delta E_{RHF}&=-U\sum_{i}\rho_{i\uparrow}\rho_{i\downarrow}=-U\sum_{q}\rho_{q\uparrow}\rho_{q\downarrow}\\ &=UN\left(-\rho_{0\uparrow}\rho_{0\downarrow}+\rho_{1}^{2}\right).\end{split}$ (17)
@@ III.1 Restricted Hartree-Fock
Figure 1: The phase diagram of model (1) in the RHF approximation, that is with $H_{t}$ given by (4) and $H^{\prime}$ given by $H^{\prime}_{RHF}$ (16). The red line separates the FM and the AFM phases. On that line, the energies of the FM and the AFM phases are equal. The calculation was done for a system on a pseudomomentum grid of size 10x10. The parameters $(\epsilon,n)$ were taken from a grid with steps $\delta\epsilon=0.03$ eV and $\delta n=1/600\approx 0.0017$.
In order to establish the energy landscape of the FM and AFM phases in the tight-binding model of the infinite-layer nickelate NdNiO2 described above, we applied the RHF method. Namely, we solved the model (1) with $H_{t}$ given by (4) and $H^{\prime}$ given by $H^{\prime}_{RHF}$ (16). There were three self-consistently determined parameters, $\rho_{0\uparrow}$ $\rho_{0\downarrow}$ and $\rho_{1}$, determined using the standard iterative procedure
$\begin{split}\rho_{0\sigma}^{l}&=\left(1-x\right)\rho_{0\sigma}^{l-1}+x\frac{1}{N}\sum_{j}\langle n_{j,\sigma}\rangle_{l-1}\\ \rho_{1}^{l}&=\left(1-x\right)\rho_{1}^{l-1}+x\frac{1}{N}\sum_{j}e^{i(\pi,\pi)\cdot\mathbf{r}(j,0)}\left\langle n_{j,\uparrow}\right\rangle_{l-1}.\end{split}$ (18)
@@
The model was defined on a 10x10 pseudomomentum lattice. In order to construct the phase diagram, we performed a scan over two free parameters. The first free parameter was the total filling $n$. The second parameter was the nickel-neodymium charge transfer energy $\epsilon$ [see Eq. (30) in Appendix A]. We investigated the parameter range
@@
with steps $\delta\epsilon=0.03$ eV and $\delta n=0.01$. This range was chosen because it covers the physical point ($\epsilon=2$ eV, $n=1$[32], as well as electron doped systems. It also covers a reasonably wide range of values for the $\epsilon$ parameter, whose value is uncertain. At each point $(\epsilon,n)$ within the range (21) we used the iterative procedure to self-consistently determine the parameters $\rho_{0\uparrow}$, $\rho_{0\downarrow}$ (FM) or $\rho_{0}$, $\rho_{1}$ (AFM). The calculation was considered to converge if
@@ III.2 Unrestricted Hartree-Fock
Figure 2: Panels (a)-(c): The value of the order parameter $X$ (see text) as a function of nickel-neodymium charge-transfer energy $\epsilon$ and total filling $n$. Results for the model (1) with kinetic energy $H_{t}$ given by (3) and Hubbard $H^{\prime}$ given by $H^{\prime}_{UHF}$ (9). Results are shown for three different values of the nickel-neodymium hopping scaling parameter: (a) $\gamma=0$, (b) $\gamma=1$, (c) $\gamma=2$. The insets show real-space maps of $S_{z}$ corresponding to particular values of $X$. The inset marked ‘Wu et. al.’ corresponds to $\epsilon$ and $n$ values used in [32]. The red line is the phase boundary, established using the RHF calculation (see text). The gray boxes indicate exotic (neither FM, nor AFM, nor paramagnetic) order. White boxes represent unconverged calculations. Panel (d): The filling on the nickel layer as a function of nickel-neodymium charge-transfer energy $\epsilon$ and total filling $n$, for $\gamma=1$. The red line visible in the plot is the same line as in panels (a)-(c).
The UHF calculations were performed as follows. We used the Hamiltonian (1) with $H_{t}$ given by (3) and $H^{\prime}$ given by $H^{\prime}_{UHF}$ (9) and found its ground state using the standard iterative procedure. The self-consistently determined parameters were $\rho_{j\uparrow}$, $\rho_{j\downarrow}$ and $\Delta_{j}$, with $j$ being the unit cell index. The iterative step was
$\begin{split}\rho_{j\sigma}^{l}&=\left(1-x\right)\rho_{j\sigma}^{l-1}+x\langle n_{j,\sigma}\rangle_{l-1}\\ \Delta_{j}^{l}&=\left(1-x\right)\Delta_{j}^{l-1}+x\left\langle c^{\dagger}_{j\uparrow}c_{j\downarrow}\right\rangle_{l-1},\end{split}$ (24)
with $x=0.2$.
@@
During the calculation, we set an upper limit on the number of iterations, the same for each point $(\epsilon,n)$ and equal to 150000. If the calculation did not converge with prescribed accuracy in this number of iterations, we marked the point as not converged. The desired accuracy was
$\max_{j\sigma}\left(\rho_{j\sigma}^{l+1}-\rho_{j\sigma}^{l},\Delta_{j}^{l+1}-\Delta_{j}^{l}\right)<10^{-6},$ (26)
where $\rho_{j\sigma}^{l}$ and $\Delta_{i}^{l}$ indicate the respective parameter values at the $l$-th step of the iterative procedure.
@@
where $f_{0}$ is the zeroth Fourier coefficient of the discrete Fourier transform of the distribution of the observable $S_{jz}$ on the 10x10 lattice
$\begin{split}S_{qz}&=\frac{1}{N}\sum_{j}e^{i\mathbf{q}\cdot\mathbf{j}}S_{jz}\\ &=f_{0}\delta(\mathbf{q})+f_{\rm AFM}\delta\left(\mathbf{q}-(\pi,\pi)\right)+\cdots,\end{split}$ (28)
where
$S_{jz}=\frac{1}{2}\left(\rho_{j\uparrow}-\rho_{j\downarrow}\right),$ (29)
@@
In the vicinity of the phase boundary, already present in the RHF calculations, non-trivial phases stabilise in all panels (a)-(c), that is for all values of $\gamma$. There are also some points for which the calculation did not converge in the maximum number of iterations  [^1] By running multiple calculations, we have verified that, although at certain points the calculation takes longer to converge, it ultimately converges to either a FM, an AFM or a more exotic phase.. Looking at Fig. 2 (a) vs (b) vs (c), one sees that the region of phase coexistence is narrowed with increasing $\gamma$, progressively replaced by the FM region. This suggests that the inter-layer coupling tends to stabilise the FM phase.

### v2

@@ Abstract
We study magnetism in the electron-doped infinite-layer nickelate NdNiO2. We perform an unrestricted Hartree-Fock calculation for a tight-binding model which contains both nickel and neodymium orbitals. We reproduce the self-doping effect, which is the escape of charge onto the neodymium bands. By fixing all free parameters to realistic values we find that undoped NdNiO2 lies right outside the antiferromagnetic (AFM) region of the phase diagram. This is consistent with experiments, which find no long-range order in the ground state of NdNiO2, yet see short-range AFM correlations and broad magnetic excitations. We also find that the self-doping effect leads to a dramatic increase in the stability of the AFM solution upon electron doping – a behavior that is strikingly different from what is, for instance, observed in the cuprates. Finally, for smaller charge transfer energies than suggested for NdNiO2, the self-doping effect may be quite strong and stabilise various stripe configurations already on the mean-field level.
@@ I Introduction
Figure 1: Visual representation of the cubic crystal structure of NdNiO2, showing the Ni ($d_{x^{2}-y^{2}}$) in whitish and Nd ($d_{z^{2}}$, $d_{xy}$) in yellow and orange colours respectively. Oxygen atoms are shown in red.
@@ II Model and Methods
$H_{t}=\sum_{i_{x},i_{y}}\sum_{i_{x}^{\prime},i_{y}^{\prime}}|i_{x},i_{y}\rangle\langle i_{x}^{\prime},i_{y}^{\prime}|\otimes h_{i_{x}-i_{x}^{\prime},i_{y}-i_{y}^{\prime}}\otimes\sigma_{0}$ (3)
describes hopping between unit cells at positions $(i_{x},i_{y})$ and $(i_{x}^{\prime},i_{y}^{\prime})$. Matrices $h$, each of dimension three, describe hopping elements in the orbital sector and $\sigma_{0}$, identity in the spin sector. In reciprocal space this becomes
$H_{t}=\sum_{\mathbf{k}}\sum_{\mathbf{d}}e^{i\mathbf{k}\cdot\mathbf{d}}|k_{x},k_{y}\rangle\langle k_{x},k_{y}|\otimes h_{d_{x},d_{y}}\otimes\sigma_{0},$ (4)
The kinetic energy is quite rich, for it follows a realistic tight-binding model with hopping elements up to third neighbor [32] and therefore we need nine matrices $h$ to describe it. We provide those in Appendix A, Eq. (37). Entries of all $h$’s are given in eV and follow [32]; matrices not given in (37) are zero.
@@
$\mathcal{H}^{\prime}=U\sum_{i_{x},i_{y}}n_{(i_{x},i_{y})3\uparrow}n_{(i_{x},i_{y})3\downarrow},$ (5)
@@
$n_{(i_{x},i_{y})3\sigma}=c^{\dagger}_{(i_{x},i_{y})3\sigma}c_{(i_{x},i_{y})3\sigma}.$ (6)
Here operator $c^{\dagger}_{(i_{x},i_{y})3\sigma}$ creates an electrons in the unit cell at point $(i_{x},i_{y})$ in orbital labeled as $3$ (or $d_{x^{2}-y^{2}}$) with spin $\sigma$. We treat $\mathcal{H}^{\prime}$ (5) in two ways. We use unrestricted Hartree-Fock (UHF) and restricted Hartree-Fock (RHF). We assume $U=4.5$ eV. Note that this value gives the $U/t\approx 12$ (where $t$ is the Ni-Ni nearest neighbor hopping element), basically in agreement with the recently postulated ratio used to describe the RIXS experiments on nickelates [33].
@@ II.1 Unrestricted Hartree-Fock approximation
In the above, the first term of the product is the projector onto the unit cell at $(i_{x},i_{y})$, the next one is the projector onto the nickel $d_{x^{2}-y^{2}}$ orbital and $\sigma_{\alpha},\;\alpha=x,y,z$ are the standard Pauli matrices. Furthermore,
$\begin{array}[]{c}\rho_{(i_{x},i_{y})\sigma}=\langle n_{(i_{x},i_{y})\sigma}\rangle\\ \Delta_{(i_{x},i_{y})}=\langle c^{\dagger}_{(i_{x},i_{y})\uparrow}c_{(i_{x},i_{y})\downarrow}\rangle,\end{array}$ (10)
where $\langle\cdot\rangle$ indicates the ground-state expectation value. The constant energy shift is
$\Delta E_{UHF}=U\sum_{i_{x},i_{y}}\left(-\rho_{(i_{x},i_{y})\uparrow}\rho_{(i_{x},i_{y})\downarrow}+|\Delta_{(i_{x},i_{y})}|^{2}\right).$ (11)
@@
In the RHF approach, we assume $\Delta_{(i_{x},i_{y})}=0$. We plug this into (8) and go to reciprocal space
$\begin{array}[]{c}\rho_{\mathbf{k}\sigma}=\frac{1}{N}\sum_{\mathbf{r}}e^{i\mathbf{k}\cdot\mathbf{r}}\rho_{\mathbf{r}\sigma}\\ c_{\mathbf{k}\sigma}=\frac{1}{\sqrt{N}}\sum_{\mathbf{r}}e^{i\mathbf{k}\cdot\mathbf{r}}c_{\mathbf{r}\sigma},\end{array}$ (12)
Furthermore, we put restriction on $\rho_{\mathbf{k}\sigma}$, limiting ourselves to ferromagnetic (FM) and antiferromagnetic (AFM) instabilities, namely
$\begin{array}[]{c}\rho_{\mathbf{k}\uparrow}=\rho_{0\uparrow}\delta(\mathbf{k})+\rho_{1}\delta(\mathbf{k}-(\pi,\pi))\\ \rho_{\mathbf{k}\downarrow}=\rho_{0\downarrow}\delta(\mathbf{k})-\rho_{1}\delta(\mathbf{k}-(\pi,\pi)).\end{array}$ (13)
@@
$\begin{split}H^{\prime}_{RHF}&=U\sum_{\mathbf{k}}\Big\{|k_{x},k_{y}\rangle\langle k_{x},k_{y}|\otimes|3\rangle\langle 3|\\ &\otimes\left[\frac{1}{2}\rho_{0\downarrow}\left(\sigma_{0}+\sigma_{z}\right)+\frac{1}{2}\rho_{0\uparrow}\left(\sigma_{0}-\sigma_{z}\right)\right]\\ &+|k_{x},k_{y}\rangle\langle k_{x}\!+\!\pi,k_{y}\!+\!\pi|\otimes|3\rangle\langle 3|\otimes\rho_{1}\sigma_{z}\Big\}.\end{split}$ (16)
@@
$\begin{split}\Delta E_{RHF}&=-U\sum_{\mathbf{r}}\rho_{\mathbf{r}\uparrow}\rho_{\mathbf{r}\downarrow}=-U\sum_{\mathbf{k}}\rho_{\mathbf{k}\uparrow}\rho_{\mathbf{k}\downarrow}\\ &=UN\left(-\rho_{0\uparrow}\rho_{0\downarrow}+\rho_{1}^{2}\right).\end{split}$ (17)
II.3 The modified Broyden Mixing
Solving self-consistent equations with many mean fields, which is the case for UHF approach, is a computationally demanding and often slow process. To deal with complications, Broyden’s method was modified to make better convergence properties [35]. This method uses $M$ previous iterations of the self-consistency loop to update the approximate inverse Jacobian. The full expression for this method is the following:
$\bm{V}^{m+1}=\bm{V}^{m}+x\bm{F}^{m}-\sum_{n=1}^{m-1}\sum_{k=1}^{m-1}\left(\omega_{n}\omega_{k}c_{k}^{m}\beta_{k,n}^{m}\bm{U}^{n}\right).$ (18)
Here $\bm{V}^{m}$ is a vector of input mean fields, i.e. the mean fields that are fed into the Hamiltonian, at $m$-th iteration and $\bm{F}^{m}$, being the residual vector, is the difference between the recalculated mean fields, after the Hamiltonian is diagonalized, and the input mean fields $\bm{V}^{m}$. Reaching self-consistency means that $\bm{F}^{m}=0$. The first part of Eq. (18) represents simple linear update where $x$ is a mixing parameter. The next part is the Broyden’s mixing correction, which ensure that the calculations are going in the right direction in the phase space. The matrix elements:
${c}^{m}_{k}=(\Delta\bm{F}^{k})^{\dagger}\bm{F}^{m},$ (19)
quantify the overlap between the current residual vector and the previous one and
$\bm{U}^{n}=x\Delta\bm{F}^{n}+\Delta\bm{V}^{m},$ (20)
is a secant update vector. We also need:
$\beta_{k,n}^{m}=[(w_{0}^{2}\bm{1}+A^{m})^{-1}]_{k,n},$ (21)
and
$A^{m}_{k,n}=w_{k}w_{n}(\Delta\bm{F}^{n})^{\dagger}\Delta\bm{F}^{k}.$ (22)
The matrix $A^{m}_{k,n}$ measures pairwise overlaps between the normalized residual vectors, while $\beta_{k,n}^{m}$ provides the corresponding weights for the update direction. The regularization term $w_{0}^{2}\bm{1}$ ensures numerical stability by preventing singularities in the matrix inverse operation. The final definitions are:
$\Delta\bm{F}^{n}=\frac{\bm{F}^{n+1}-\bm{F}^{n}}{|\bm{F}^{n+1}-\bm{F}^{n}|},$ (23)
$\Delta\bm{V}^{m}=\frac{\bm{V}^{m+1}-\bm{V}^{m}}{|\bm{F}^{m+1}-\bm{F}^{m}|}.$ (24)
Symbol $\bm{1}$ in equation for $\beta^{m}$ denotes $(m-1)\times(m-1)$ identity matrix. The weights $w_{k},w_{n}$ are usually equal to 1. The regularization parameter $w_{0}$ is usually fixed to 0.01. In our implementation it is adaptively updated at each iteration according to the condition number of $A^{m}_{k,n}$ in order to avoid instabilities and improve convergence robustness.
Note that the correction part of the Broyden mixing can use a finite number of previous iterations. We choose this number to be 100 which means that after 100 iterations the history is reset and rebuilt. This choice reduces sensitivity to the random initialization, keeps the cost of calculation controlled, and prevents the mixer from overfitting stale history – issues that can stall convergence to an undesired fixed point. Restarting the history can also help the algorithm escape metastable solutions.
@@ III.1 Restricted Hartree-Fock
Figure 2: The phase diagram of model (1) in the RHF approximation, that is with $H_{t}$ given by (4) and $H^{\prime}$ given by $H^{\prime}_{RHF}$ (16). The red line separates the FM and the AFM phases. On that line, the energies of the FM and the AFM phases are equal. The calculation was done for a system on a pseudomomentum grid of size $10\times 10$. The parameters $(\epsilon,n)$ were taken from a grid with steps $\delta\epsilon=0.03$ eV and $\delta n=1/600\approx 0.0017$.
In order to establish the energy landscape of the FM and AFM phases in the tight-binding model of the infinite-layer nickelate NdNiO2 described above, we applied the RHF method. Namely, we solved the model (1) with $H_{t}$ given by (4) and $H^{\prime}$ given by $H^{\prime}_{RHF}$ (16). There were three self-consistently determined parameters, $\rho_{0\uparrow}$ $\rho_{0\downarrow}$ and $\rho_{1}$, determined using the standard iterative procedure with linear mixing
$\begin{split}\rho_{0\sigma}^{l}&=\left(1-x\right)\rho_{0\sigma}^{l-1}+x\frac{1}{N}\sum_{\mathbf{r}}\langle n_{\mathbf{r},\sigma}\rangle_{l-1}\\ \rho_{1}^{l}&=\left(1-x\right)\rho_{1}^{l-1}+x\frac{1}{N}\sum_{\mathbf{r}}e^{i(\pi,\pi)\cdot\mathbf{r}}\left\langle n_{\mathbf{r},\uparrow}\right\rangle_{l-1}.\end{split}$ (25)
@@
The model was defined on a $10\times 10$ pseudomomentum lattice. In order to construct the phase diagram, we performed a scan over two free parameters. The first free parameter was the total filling $n$. The second parameter was the nickel-neodymium charge transfer energy $\epsilon$ [see Eq. (37) in Appendix A]. We investigated the parameter range
@@
with steps $\delta\epsilon=0.03$ eV and $\delta n=0.01$. This range was chosen because it covers the physical point ($\epsilon=2$ eV, n=1n=1[32], as well as electron doped systems. It also covers a reasonably wide range of values for the $\epsilon$ parameter, whose value is uncertain. At each point $(\epsilon,n)$ within the range (28) we used the iterative procedure to self-consistently determine the parameters $\rho_{0\uparrow}$, $\rho_{0\downarrow}$ (FM) or $\rho_{0}$, $\rho_{1}$ (AFM). The calculation was considered to converge if
@@ III.2 Unrestricted Hartree-Fock
Figure 3: Panels (a)-(c): The value of the order parameter $X$ (see text) as a function of nickel-neodymium charge-transfer energy $\epsilon$ and total filling $n$. Results for the model (1) with kinetic energy $H_{t}$ given by (3) and Hubbard $H^{\prime}$ given by $H^{\prime}_{UHF}$ (9). Results are shown for three different values of the nickel-neodymium hopping scaling parameter: (a) $\gamma=0$, (b) $\gamma=1$, (c) $\gamma=2$. The insets show real-space maps of $S_{z}$ corresponding to particular values of $X$. The inset marked ‘Wu et. al.’ corresponds to $\epsilon$ and $n$ values used in [32]. The red line is the phase boundary, established using the RHF calculation (see text). The gray boxes indicate exotic (neither FM, nor AFM, nor paramagnetic) order. Panel (d): The filling on the nickel layer as a function of nickel-neodymium charge-transfer energy $\epsilon$ and total filling $n$, for $\gamma=1$. The red line visible in the plot is the same line as in panels (a)-(c).
The UHF calculations were performed as follows. We used the Hamiltonian (1) with $H_{t}$ given by (3) and $H^{\prime}$ given by $H^{\prime}_{UHF}$ (9) and found its ground state using the standard iterative procedure. The self-consistently determined parameters were $\rho_{\mathbf{r}\uparrow}$, $\rho_{\mathbf{r}\downarrow}$ and $\Delta_{\mathbf{r}}$, with $\mathbf{r}=(i_{x},i_{y})$ being the unit cell index. The iterative step was
$\begin{split}\rho_{\mathbf{r}\sigma}^{l}&=\left(1-x\right)\rho_{\mathbf{r}\sigma}^{l-1}+x\langle n_{\mathbf{r},\sigma}\rangle_{l-1}\\ \Delta_{\mathbf{r}}^{l}&=\left(1-x\right)\Delta_{\mathbf{r}}^{l-1}+x\left\langle c^{\dagger}_{\mathbf{r}\uparrow}c_{\mathbf{r}\downarrow}\right\rangle_{l-1},\end{split}$ (31)
with linear mixing parameter $x=0.2$.
@@
During the calculation, we set an upper limit on the number of iterations, the same for each point $(\epsilon,n)$ and equal to 150000. The desired accuracy was
$\max_{\mathbf{r}\sigma}\left(\rho_{\mathbf{r}\sigma}^{l+1}-\rho_{\mathbf{r}\sigma}^{l},\Delta_{\mathbf{r}}^{l+1}-\Delta_{\mathbf{r}}^{l}\right)<10^{-6},$ (33)
where $\rho_{j\sigma}^{l}$ and $\Delta_{i}^{l}$ indicate the respective parameter values at the $l$-th step of the iterative procedure. If for a given point in the phase diagram the calculation did not converge with prescribed accuracy within iterations limit, we used modified Broyden mixing for this point.
@@
where $f_{0}$ is the zeroth Fourier coefficient of the discrete Fourier transform of the distribution of the observable $S_{\mathbf{r}z}$ on the $10\times 10$ lattice,
$\begin{split}S_{\mathbf{k}z}&=\frac{1}{N}\sum_{j}e^{i\mathbf{k}\cdot\mathbf{r}}S_{jz}\\ &=f_{0}\delta(\mathbf{k})+f_{\rm AFM}\delta\left(\mathbf{k}-(\pi,\pi)\right)+\cdots,\end{split}$ (35)
with
$S_{\mathbf{r}z}=\frac{1}{2}\left(\rho_{\mathbf{r}\uparrow}-\rho_{\mathbf{r}\downarrow}\right),$ (36)
@@
In the vicinity of the phase boundary, already present in the RHF calculations, non-trivial phases stabilise in all panels (a)-(c), that is for all values of $\gamma$. Looking at Fig. 3 (a) vs (b) vs (c), one sees that the region of phase coexistence is narrowed with increasing $\gamma$, progressively replaced by the FM region. This suggests that the inter-layer coupling tends to stabilise the FM phase.
@@
III.3 Unrestricted Hartree-Fock with the modified Broyden mixing
As the final step, we replace the linear mixing by the modified Broyden mixing in the UHF scheme. Using this method we perform calculations for the same set of parameters as in subsection  III.2 – except that all the calculations are done solely for the realistic nickel-neodymium scaling parameter $\gamma=1$. The obtained phase diagram can be seen in Fig. 4.
Figure 4: The value of the order parameter $X$ (see text) as a function of nickel-neodymium charge-transfer energy $\epsilon$ and total filling $n$ calculated for the nickel-neodymium scaling parameter $\gamma=1$ using the UHF approximation with the modified Broyden mixing. The red line is the phase boundary, established using the RHF calculation (see text).
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 5: Spin configurations of the non-collinear orders, stable in the unrestricted Hartree-Fock calculations with the modified Broyden mixing of Sec. III.3: ($\blacktriangle$): Spin configuration for the AFM axial stripe order stable for charge transfer energy $\epsilon=0.5$ eV and total filling $n=1.15$. ($\blacksquare$): Spin configuration for the FM axial stripe stable for $\epsilon=0.5$ eV and $n=1.15$. ($\bullet$): Diagonal stipe spin configuration stable for $\epsilon=1.1$ eV and $n=1.4$. ($\star$): Vortex spin configuration stable for for $\epsilon=1.1$ eV and $n=1.15$.
The results show that the AFM region is essentially unchanged relative to the results of the previous section, cf. Fig. 4 vs. Fig. 3. On the other hand, the phase space taken by the FM order diminishes. Instead various noncollinear phases set in. Interestingly, the ‘standard’ UHF approach from the previous section only supports collinear phases.
[…]
Interestingly, the diagonal stripes (green in Fig. 4) and the axial stripes (magenta/red in Fig. 4) occur in roughly equal proportions, showing no strong orientation preference. Fig. 5 visualizes the three stable stripe and one incommensurate texture.